### v1

# Designing Speech Technologies for Australian Aboriginal English: Opportunities, Risks and Participation

Abstract.
In Australia, post-contact language varieties, including creoles and local varieties of international languages, emerged as a result of forced contact between Indigenous communities and English speakers. These contact varieties are widely used, yet are poorly supported by language technologies. This gap presents barriers to participation in civil and economic society for Indigenous communities using these varieties, and reproduces minoritisation of contemporary Indigenous sociolinguistic identities. This paper concerns three questions regarding this context. First, can speech technologies support speakers of Australian Aboriginal English, a local indigenised variety of English? Second, what risks are inherent in such a project? Third, what technology development practices are appropriate for this context, and how can researchers integrate meaningful community participation in order to mitigate risks? We argue that opportunities do exist—as well as risks—and demonstrate this through a case study exploring design practices in a real-world project aiming to improve speech technologies for Australian Aboriginal English. We discuss how we integrated culturally appropriate and participatory processes throughout the project. We call for increased support for languages used by Indigenous communities, including contact varieties, which provide practical economic and socio-cultural benefits, provided that participatory and culturally safe practices are enacted.
1. Introduction
Most work on speech and language technologies, a.k.a. Natural Language Processing (NLP), focuses on languages which are spoken by majority populations and heavily standardised within one or more countries (Bender, 2009; Sawicki et al., 2023; Joshi et al., 2020). The field of NLP is thus reproducing the minoritisation of language varieties, and in the process reinforcing barriers of exclusion for diverse language communities. These include most local languages in highly linguistically diverse countries such as Indonesia (Aji et al., 2022), as well as various creoles which combine elements of settler-colonial languages elsewhere in the world (Lent et al., 2021). Additionally affected are linguistic varieties of national languages spoken by minoritised groups (Kantharuban et al., 2023), such as Māori English in New Zealand, Multicultural London English in the UK, and African-American English in the US. Speakers of minoritised varieties must typically switch to a majority variety when using language technologies, and when they do they can experience negative psychological impacts, such as feeling that technology wasn’t designed for them (Cunningham et al., 2024; Mengesha et al., 2021; Harrington et al., 2022; Wenzel et al., 2023).
Machine learning-driven NLP is not typically oriented towards participatory processes, driven instead by the valorisation of efficiency (Birhane et al., 2022b) and fundamentally constrained by tensions between participation and scale (Young et al., 2024). NLP evaluations emphasise quantitative and decontextualised metrics (Hutchinson et al., 2022), while language dataset projects often reduce participation to crowdsourced platform work. These projects fail to create relationships which emphasise reciprocity and social justice (Sloane et al., 2022; Cooper and Zafiroglu, 2024). Participatory AI projects, on the other hand, are challenged by a lack of shared understanding of good practices (Groves et al., 2023), and may risk “participation-washing” (Sloane et al., 2022). More generally, projects which aspire to participation often fail to afford decision-making power to participating communities (Arnstein, 1969), and risk co-optation when delivered through corporate structures (Asaro, 2000; Birhane et al., 2022a).
In Australia, where this research takes place, of the 400+ ancestral Indigenous languages that existed before European colonisation (Bowern, 2023, p. 56), most are either extinct or critically endangered (Australian Government et al., 2020). This is due to linguicidal practices of colonialism, which at various times included governments forcibly separating Aboriginal children from their families and coercing Aboriginal pupils to speak English while forbidding them from speaking their languages (Rademaker, 2018). As a result, most Indigenous people in Australia speak Aboriginal English, making it a strong encoder of ethno-cultural identity (Rodríguez Louro and Collard, 2024). This indigenised variety of English integrates phonological, lexico-grammatical and discourse-pragmatic features from English and Aboriginal ancestral languages (see §2.1).
Given this context, we take three positions in this paper. First, Indigenous people should be supported to use whichever of their linguistic varieties that they prefer, including contact varieties, when using language technologies. Second, NLP requires not only culturally appropriate outputs, such as models, tasks and evaluations (Hovy and Yang, 2021; Hershcovich et al., 2022; Liu et al., 2022), but also culturally appropriate processes, outcomes and modes of participation (Noe and Kirshenbaum, 2024; Bird and Yibarbuk, 2024). All are currently undervalued in NLP. Third, language technologists should be equally concerned with contemporary Indigenous needs, including supporting recent and evolving Indigenous language varieties, as they are with needs related to ancestral languages. What, then, are these opportunities, as well as the risks? And how are they distinct from the technology needs of mainstream language communities, or of ancestral Indigenous language communities?
To address these questions, this paper contributes the first detailed consideration of language technologies to support Aboriginal English speakers. To our knowledge, this is the first paper to focus on technologies for this language community. We do so through a case study of an ongoing project intended to improve Automatic Speech Recognition (ASR) for Indigenous communities in Australia. In doing so, this paper follows previous NLP literature in considering a specific Indigenous language community and its context (e.g., Zhang et al., 2022; Teodorescu et al., 2022), with a focus on design considerations in order to avoid assuming that mainstream methods and goals are locally applicable (Noe and Kirshenbaum, 2024; Pine et al., 2024). As such, this paper contributes to our understanding of ethical and responsible development of technologies for minoritised communities, as well as to our understanding of how to design and build culturally sensitive technologies while incorporating participatory methods.
The paper begins with an introduction to Australian Aboriginal English, including a description of relevant linguistic and sociolinguistic factors (§2). We also present a brief survey of related work on Indigenous language technologies. Next, we discuss the opportunities and risks inherent in deploying ASR systems for this community, as well as opportunities and risks in projects building such technologies (§3). Building on this analysis, we explore strategies to mitigate risk, and facilitate opportunities, through a detailed discussion of the design considerations of a real-world project to design culturally appropriate ASR while integrating participatory methods (§4). In doing so, we provide a detailed discussion of some of the considerations that are critical for a project of this sort, including system requirements, data elicitation and transcription, system evaluation, and processes and frameworks for participation, and employment. Finally, we discuss further technological opportunities for supporting speakers of Aboriginal English (§5).
2. Background and Related Work
When considering how technologies can support speakers of a specific Indigenous language variety, it is critical to have some basic knowledge of the speech community, the language variety, and their contexts (Zhang et al., 2022). In this section we introduce Aboriginal English and its context, followed by a brief summary of prior work relevant to language technologies for Indigenous communities. This includes consideration of both ancestral and contact language varieties, and also of participatory methods in Indigenous language projects.
2.1. Australian Aboriginal English
Aboriginal English is a contact variety of English used by Indigenous people in Australia (Rodríguez Louro and Collard, 2021). It is particularly prevalent across southern Australia where ancestral languages are less widely spoken. In these contexts, Aboriginal English is used to strengthen Indigenous ways of being and communicating (Eades, 2012a). Aboriginal English has its origins in the product of forced contact between Indigenous people and the English-speaking settler-colonialists. The English spoken by the British was a powerful instrument used to strip Indigenous communities of their ancestral rights (Malcolm, 2000, p. 125). Aboriginal English’s precursor was a mixed jargon viewed as ‘broken English’ by the British. By the time this jargon had transitioned to a novel pidgin (known as ‘New South Wales pidgin’ (Troy, 1994)), it began to travel outside present-day Sydney and across New South Wales. With the advent of the pastoral industry, the pidgin also spread north into Queensland and the Northern Territory, giving rise to Queensland Pidgin English and the Northern Australian creoles, including Roper River Kriol (Meakins, 2014, p. 371).
The Aboriginal English has linguistic characteristics, including pronunciation, lexis, grammar, and discourse-pragmatics, that distinguish it from mainstream English (Eades, 2012a). Pronunciation features include: frequent /h/-dropping and /h/-addition; metathesis of ‘k’ and ‘s’ in the word ‘ask’, which is pronounced ‘aks’; and default reduction of the ‘ing’ verbal suffix to ‘in’. Grammatical features include double negatives as in ‘never done nothing’; and double subjects as in ‘My niece, she did it.’ Lexical features include lexical transfer such as blackfella, meaning ‘black person/fellow’, and borrowings from ancestral languages, such as dardy ‘good/good looking’ from Nyungar.
Experts remain ambivalent as to whether Aboriginal English represents a single ethnolect, or whether different varieties of Aboriginal English, or ‘Aboriginal Englishes’, exist across Australia (Dickson, 2019). Aboriginal English is frequently placed on a continuum that ranges from ‘lighter’ acrolectal varieties (closer to standardised Australian English) to ‘heavier’ basilectal varieties which are closer to Kriol, spoken across northern Australia (Schultze-Berndt et al., 2013). The basilectal varieties are mostly unintelligible to standardised English speakers, but the acrolectal varieties are superficially similar to Australian English.
Because Aboriginal English is English based, it is sometimes socially stigmatised, erroneously classified by laypeople as an incorrect version of English. As a result, L1 users of Aboriginal English are often discriminated against in schooling and legal settings (Harkins, 1994; Eades, 2012b), perpetuating invisibility of the variety and leading to the false perception that its speakers are uneducated and rude (Eades, 2013). Stigmatisation of Aboriginal English is evident in the comment below, which was offered in response to an online article on the role of Aboriginal English in Indigenous health communication (Rodríguez Louro and Collard, 2021).
With Aboriginal children all going to school, and nearly all of them being taught in English, does it still serve them to support the continued use of Pidgin English?
The commenter’s deliberate use of ‘Pidgin English’ reflects a deficit perspective that undermines the value of Aboriginal English. Rather than embracing Aboriginal English as a linguistic variety many Indigenous people use daily in Australia, this perspective portrays racialised peoples’ language varieties as ‘lacking’ vis-à-vis mainstream English (Hudley et al., 2020, p. e214).
These negative attitudes are exacerbated because Aboriginal English is primarily oral, with no nationally adopted guidelines as to how it should be written. Additionally, the orthographic representation of Aboriginal English varies across locales, reflecting the local linguistic biographies and practices of distinct Indigenous groups. For example, the name of the Traditional Owners of the region comprising the southwest of Western Australia, as well as their language, is variably spelled ‘Nyungar’, ‘Noongar’, or ‘Nyoongar’ depending on community members’ preferences. [^1] Author 3 is Nyungar so we prefer this spelling. Rather than a weakness, the lack of standardising orthographies for Aboriginal English reflects Indigenous people’s prerogative to differentiate—rather than homogenise—how Aboriginal English is used across Indigenous Country.
Indigenous people tend to find the use of direct questions awkward or offensive (Rodríguez Louro and Collard, 2024), [^2] With some exceptions, including teaching contexts where questions instructors asking questions is heavily routinised (Moses et al., 2008). and these pragmatic considerations continue in Aboriginal English. Direct questions are also likely to elicit misleading information; Indigenous people have been known to answer ‘yes’ or agree with non-Indigenous people even if the question is not fully understood—a phenomenon known as ‘gratuitous concurrence’ (Liberman, 1985, p. 137). To avoid these limitations, researchers have used ‘yarning’, an Indigenous cultural form of storytelling and conversation, to lead unstructured data collection with Indigenous communities who speak Aboriginal English (Anonymous, YYYY).
2.2. Technologies for ancestral Indigenous languages
The NLP literature for Indigenous languages tends to focus on ancestral languages, i.e., Indigenous languages which predate colonisation. [^3] Others call these languages ‘traditional’ Indigenous languages, e.g., (Eades, 2013). The term ‘ancestral’ is sometimes used by others with a different meaning: a language from which modern languages descend. Given the considerable overlap between endangered languages and Indigenous languages (Bromham et al., 2022)—a consequence of colonisation—NLP literature tends to focus on application domains which support language revitalisation and preservation efforts, including assisting language learners or teachers (Meighan, 2021, 2022; Galla, 2009). This literature discusses technologies such as quiz generation, automated language assessment, language learning platforms, computer assisted translation, predictive text, and readability classification (e.g. Zhang et al., 2022; Teodorescu et al., 2022; Carew et al., 2015; Mager et al., 2024). Bird and Yibarbuk (2024) pose open questions regarding distinct needs of learners of oral languages without the technology working through a written form.
Language preservation or documentation is frequently cited as a goal (Le Ferrand, 2023; Prud’hommeaux et al., 2021), and is discussed variously as a means towards revitalisation, or as an end in itself. Recognising that ancestral language communities can be minoritised in multiple ways, Carew et al. (2015) discuss the role of language technologies in promoting digital inclusion. Other motivations encompass the importance of language technologies in cultural revitalisation projects, e.g., helping to learn about traditional place names, stories, songs and kinship relations.
2.3. Technologies for contact language varieties
The NLP literature has to-date had marginal concern with supporting Indigenous contact language communities, such as Hawaiian Pidgin, Australian Kriol, Tok Pisin, Chavacano, etc. For example, Nigerian Pidgin has over 100 million speakers (Lin et al., 2024), yet is poorly supported by technology. This seems to be due to a combination of reasons, including the stigmatisation of contact languages (Siegel, 1999; Lent et al., 2024) and a belief that ancestral languages are more important in terms of global heritage. However Indigenous peoples are not museum artifacts or relics of an ‘ancient’ era (cf.  Gaby and Woods, 2020, e274); their cultures and societies are dynamic and evolving.
To understand the needs of Indigenous contact language communities, we first draw on the NLP literature for Creole languages. Lent et al.’s survey of six Creole communities found different levels of desire for language technologies (Lent et al., 2022). They note that these languages often lack recognised status, are primarily oral with variable orthography and writing conventions, and are often socially stigmatised. This points to the importance of contact varieties in shaping contemporary social identities. If technology does not support contemporary forms of Indigenous expression, then it reproduces and reinforces minoritisation of Indigenous culture and identity. Contact varieties are often vehicular, used to communicate with other communities, including participation in national and regional economies, and with civic services (Bird, 2022). If language technologies are to support economic and civic participation, then they need to support vehicular varieties.
Since contact varieties are primarily oral, prior work has identified speech technologies as the most wanted by both experts and communities (Lent et al., 2022; Le Ferrand, 2023). Robust Automated Speech Recognition (ASR) that can recognise Indigenous pronunciations and loan words is an important foundational technology for a broad range of applications, from voice search to automated transcription and interaction with digital assistants (Aksënova et al., 2021). Design of ASR for contact varieties should be mindful of the social dimensions of orthography, in particular of variation from standardised written forms.
For some contact varieties, orthographic norms may have colonial associations, and can be alienating for speakers (Markl et al., 2023). For this reason, and due to their primary orality, consultation with experts and communities is required to ascertain the need for text-based technologies (Lent et al., 2022; Le Ferrand et al., 2022; Bird, 2022). When needed, writing tools supporting local scripts, fonts, and keyboards would generally be the most useful text-based technologies (Esch et al., 2019). The lack of standardised writing conventions for many contact varieties may present challenges for the utility of technologies such as spell checkers and next word predictors.
Language technologies are increasingly multilingual, and a critical component of many systems is Language Identification (LangId). Since most Indigenous communities are multilingual, robust LangId and code-switching/mixing detection are needed to support local multilingual technologies (Nguyen and Dogruöz, 2013).
2.4. Culture, and participatory methods, in language research
Recently, the NLP community has demonstrated increasing interest in the topic of culture, with, e.g., two recent workshops on Cross-cultural Considerations in NLP (Dev et al., 2023; Prabhakaran et al., 2024). This interest has been spurred by the capabilities of Large Language Models (LLMs) with regards to linguistic form, while also demonstrating lack of cultural nuance (Hershcovich et al., 2022; Naous et al., 2024; Li et al., 2024). Threads of research have focused on measuring cultural biases (Naous et al., 2024; Tao et al., 2024) and expression of cultural values by LLMs (Arora et al., 2023; Durmus et al., 2023; Ramezani and Xu, 2023).
It is critical that the NLP community goes further than developing technological fixes in data and models, to consider design questions and project processes. NLP research using Indigenous language data often fails to engage with the cultural differences between technologists and the communities they aspire to support (Bird and Yibarbuk, 2024). Recent community-led projects such as Masakhane (Nekoto et al., 2020) and the Papa Reo project (Jones et al., 2023) provide alternative models for the NLP community, by drawing on principles of participatory design to embed the cultural knowledge of language communities throughout NLP research and development processes. Rather than developing technologies for Indigenous contact varieties in isolation of communities speaking those varieties, the primary focus of such projects must be on developing language technologies with Indigenous language communities. This requires research teams, and the broader NLP community, to consider the position of NLP researchers relative to community partners, reflecting not only on performance disparities of technologies across languages or cultural biases, but also on the composition of research and development teams, as the cultural assumptions and norms of such teams influence the research we conduct (Forsythe, 2001).
Within sociolinguistics, the ‘yarning’ method has demonstrated the effectiveness of culturally appropriate, participatory research methods when working with Aboriginal communities on speech data collection projects (Anonymous, YYYY). ‘Yarning’ favours group-based unstructured storytelling over individual interviews, which aligns with the Indigenous cultural protocol of centring the group over the individual (Malcolm, 2018, p. 17). Engaging in groups makes people feel less self-conscious in speech data collection activities than in one-on-one interactions with a researcher. Previous research led by Author 3, who is Indigenous, demonstrates how using such culturally appropriate, participatory methods throughout the research process—from choosing data collection protocols and fieldwork location to participant choice—enables recording of the authentic voices of people infrequently represented in academic research.
2.5. Projects with Indigenous communities
The literature emphasises that working with data of Indigenous language varieties raises ethical considerations. These also apply to contact varieties such as Australian Aboriginal English, and are critical to keep in mind when considering opportunities and risks. Four questions that recur in the literature include:
Q1. What is the context?
Before working with Indigenous varieties, researchers must reflect on the harmful impact of past colonial research projects on Indigenous communities and seek to understand the historical and sociolinguistic context of partner communities (Smith, 2021; Schwartz, 2022; Doğruöz and Sitaram, 2022; Bird, 2020a; Zhang et al., 2022; Noe and Kirshenbaum, 2024; Le Ferrand et al., 2022).
Q2. What relationships will be maintained?
Relationships are fundamental to Indigenous identities, communities, and research (Smith, 2021; Wilson, 2020), and this holds for research on contact varieties. Technologists should be cognisant that relationships with partners and participants may, and in some cases should, persist beyond the scope of the project (Cooper et al., 2024).
Q3. Who will control the project, and its outputs?
In both ancestral and contact variety-speaking communities, Indigenous people are less likely to instigate language technology projects, enumerate technological goals, or design/control project budgets, and it is important to both acknowledge and address these power differentials. Using language data ethically is the lowest threshold for language projects (Cameron et al., 1993), and involves considerations not only of IP (Intellectual Property) but also ICIP (Indigenous Cultural and Intellectual Property) (Janke, 2019).
Q4. Is the technology appropriate for the community?
Indigenous varieties are typically differentiated functionally within a multilingual community, and technologies may be more useful for some contact varieties than others (Doğruöz and Sitaram, 2022; Bird, 2022; Lent et al., 2022). Technologists should understand the technology needs of Indigenous communities, which may be distinct from those of non-Indigenous users (Liu et al., 2022; Cooper et al., 2024).
3. Opportunities and Risks for ASR for Australian Aboriginal English
With the contexts and questions of the previous section in mind, we now consider Automated Speech Recognition (ASR) systems for Aboriginal English, and the potential opportunities and risks. We consider both deployed ASR systems as well as the projects which build ASR systems; our ultimate goal is to build inclusive technologies using inclusive methods.
3.1. System Deployment Opportunities
ASR supports many application domains (Aksënova et al., 2021), and improving ASR for Aboriginal English could provide multiple specific benefits to Aboriginal and Torres Strait Islander people, their communities, and non-Indigenous people who are encountering Aboriginal English. We attempt to distinguish those possible benefits with as much specificity as possible, and in contrast to prior work we adopt a human-centered framing rather than a technology-centered one.
Opportunity 1: Improved dictation for Aboriginal English speakers.
Dictation typically involves a single speaker consciously speaking directly to a device, and is useful in a variety of contexts in which speaking is preferred to typing. These include messaging while driving, speakers with physical or vision disabilities, speakers avoiding small keyboards on mobile devices, and speakers with dyslexia. The speaker often reviews the text output for accuracy. Since they are conscious of the interaction, some speakers may modify their speech in an attempt to improve the fidelity of the transcription, such as minimising use of informal registers, obscure or complex words, and minoritised varieties (Mengesha et al., 2021; Holt et al., 2024).
We note that not all dictation applications are likely to benefit Aboriginal English speakers. For example, ASR is sometimes used in dictation practice by second-language learners, however Aboriginal English is not formally taught or studied as a second-language.
Opportunity 2: Improved transcriptions for readers of personal messages.
ASR can be initiated by those within the speaker’s community or social network, for example, users of a voicemail service providing automated transcriptions. This can be useful in a variety of contexts in which reading is preferred to listening, including in quiet environments in which silence is preferred, in noisy environments in which hearing is difficult, when privacy is desirable in public environments, and for people with hearing disabilities. An Aboriginal English speaker is often unaware when their voice is being transcribed, and the reader might or might not listen to the voice recording to verify fidelity of the transcription. The reader may be able to accommodate and even correct some ASR errors, based on their familiarity with the speaker and/or context, even without listening to the recording.
Opportunity 3: Improved transcriptions for consumers of platform content.
When speech and video platforms host content, they often provide closed captions. Such content can be unscripted (typical for Aboriginal English on YouTube, for example, see Appendix B) or scripted (common in TV shows, movies, podcasts and vodcasts, etc—for an analysis of scripted Aboriginal English see (Bednarek, 2021)). These are generated offline and stored as metadata alongside the Aboriginal English voice recording. The platforms may have some access controls, but in the general case both the recordings and the transcriptions may be accessible to potentially anyone online, including listeners and readers with no prior experience with Aboriginal English. Listening to the recordings might not help readers without familiarity with Aboriginal English to correct some ASR errors.
Due to systemic societal biases, some applications of closed captioning may encounter relatively little Aboriginal English compared to Australian English. The automated transcription of Australian parliamentary proceedings (Gollan et al., 2005), for example, is one use case for Aboriginal English ASR that is impacted both by the under-representation of Indigenous people in Australian parliaments and by the social stigmatisation of speaking Aboriginal English in formal contexts.
Opportunity 4: Improved human-machine interactions for Aboriginal English speakers.
ASR is increasingly incorporated into a wide range of computational systems used by Aboriginal English speakers. These include automated voice assistants, automated or semi-automated call centres including for critical banking and government services, web search using voice queries, and audio/video retrieval that indexes speech under transcribed terms. (Voice Query-By-Example technologies (Parada et al., 2009) do not require ASR but constitute a minority of audio retrieval systems.) In many cases, the transcription might not be seen by any human, including the Aboriginal English speaker. Instead, the transcription is typically the input to a Natural Language Understanding component, which outputs some form of semantic representation in forms such as a task frame, a search index entry, or an expanded Information Retrieval query (Efthimiadis, 1996; Jurafsky and Martin, 2008). More accurate transcriptions can lead to more relevant search results, more efficient interactions, and fewer abandoned tasks.
3.2. System Deployment Risks
Schwartz has argued that the primary ethical responsibility when building language technologies for Indigenous communities is to not do any harm (Schwartz, 2022). Although many taxonomies of risks and negative impacts of technologies exist, our goal here is be a specific and situated as possible. Hence we do not attempt to map the identified risks to prior taxonomies which often aim at universalism and generality (Berman et al., 2024).
Risk 1: Inappropriate for the contexts of Aboriginal English.
Researchers should not assume that minoritised language communities have the same wants and needs as mainstream language communities (Bird, 2022), nor that multilingual speech communities require the same technologies for all languages (Lent et al., 2022; Bird and Yibarbuk, 2024). For example, would some Aboriginal English speakers prefer to code-switch into Standardised Australian English when interacting with machines, if it results in greater task efficiency? What about when dictating messages to family and friends? Conversely, would they prefer more support to speak Aboriginal English in contexts which are currently not well supported? There is a shortage of research on the real-world needs of Aboriginal English speakers. Building systems that are neither needed or wanted is not only inefficient—it also undermines the rights of minoritised communities to determine which technologies are built for them.
Risk 2: Lack of sociolinguistic ecological validity.
Like all language varieties, Aboriginal English is not homogeneous, but exhibits regional variation (Dickson, 2019). Furthermore, and notably, Aboriginal English includes both acrolectal and basilectal varieties (§2.1). If the training data for an ASR model does not represent sufficient variation, then the model will fail to recognise the breadth of forms of Aboriginal English. If the training data collection protocols are not designed to be culturally safe, then speakers may code switch into Standardised Australian English, reducing the validity of the data.
Risk 3: Psychological harms for Aboriginal English users when dictating or interacting with systems.
Prior research has found that when ASR systems make errors, some members of minoritised language communities feel that the computer “doesn’t like” how they are speaking (Brewer et al., 2023), or that the technology isn’t built for people like them (Mengesha et al., 2021). Whereas speakers of mainstream language varieties may blame the ASR system for poor performance, speakers of minoritised communities may blame themselves, and experience a drop in self esteem or an increase in self consciousness about their identity (Wenzel et al., 2023). They may feel that their language variety is being judged as illegitimate (Cunningham et al., 2024). If ASR fails to meet expectations for speakers of Aboriginal English, all of these impacts may be reproduced.
Risk 4: Torquing of Aboriginal English sociolinguistic identities.
ASR systems function as classifiers not just of speech, but also, by extension, of language varieties and sociolinguistic identities. As such, ASR systems can be examined through Bowker and Star’s sociological lens of torque and residuals (Bowker and Star, 2000). “Residuals” refers to the “other” categories that are relevant to the individual but missing from the classificatory system, while “torque” refers to the painful twisting and bending that occurs when an individual’s data is forced into an inappropriate classificatory system, such as forcing individuals into racial categories in Apartheid South Africa. Previous work has explored how gender classification technologies produce torque and othering for trans people (Scheuerman et al., 2019), and how medical and biometric IT projects can torque minoritised groups (Avlona and Shklovski, 2024; Singh and Jackson, 2017). We posit that a similar form of torque exists for users of non-standardised language varieties, when ASR systems systematically fail to recognise spoken forms salient to social identities, bending and twisting them into the written forms of mainstream dictionaries of colonising cultures.
3.3. Project Opportunities
We take the position that language technology projects provide the potential for a broad range of opportunities, especially when the project is conducted not just for the community but also with the community (Cameron et al., 1993).
Opportunity 5: Project leads to greater non-Indigenous understanding of Indigenous contexts and histories.
Projects involving both Indigenous and non-Indigenous participants create opportunities for ‘two way learning’, a phrase used in Australia to refer to Indigenous and non-Indigenous people coming together to learn from each other in a space that is safe for those with diverse perspectives and ways of knowing (Purdie et al., 2012, p. xx).
Opportunity 6: Project produces economic opportunities for Indigenous community.
Data collection projects can provide opportunities for what Sloane et al. (2022) call “participation as work”, with equitable compensation. There are large economic disparities in Australia, with median incomes for Indigenous people around 23% lower than for non-Indigenous (Australian Institute for Health and welfare, 2023), and Indigenous people have higher rates of unemployment for all education levels (Australian Institute of Health and Welfare, 2023b). If equitable compensation and non-extractive processes are successfully adopted, data work can create economic opportunities for community.
Opportunity 7: Project builds capabilities in community.
Building Indigenous capabilities on technology projects has the opportunity of addressing the current digital divide in skills (Carew et al., 2015). When projects effectively build community capacity, participants can draw on their developing technical skills—such as data collection and management capabilities—to advocate for their communities in future projects and on technology and policy issues (Cooper et al., 2022). To achieve these benefits, meaningful opportunities for skill development need to be incorporated into the project, focusing on capabilities that will remain valuable to the community beyond the project’s completion.
Opportunity 8: Project increases Indigenous self-determination.
Providing opportunities for Indigenous control over project outputs and design decision can have not only consequential benefits (e.g., avoiding data set misuse by third parties in ways that do not accord with Indigenous aspirations), but also constitute a form of procedural justice through greater self-determination.
3.4. Project Risks
Developing inclusive ASR inclusively requires more than addressing model biases (Feng et al., 2024)— it also requires researchers to integrate participatory practices throughout the project. As such, we take the position that, when developing language technologies, risks need to be considered using a wide-angle lens that takes into account not just the deployment scenarios but also the project processes used when developing the technology. For projects with Indigenous communities, “it’s how do you things that matters” (Cooper et al., 2024), and the ends should not be used to justify the means.
Risk 5: Project harms relationships with community.
The literature on building language technologies for Indigenous communities emphasises the foundational importance of relationships with the community (Zhang et al., 2022; Cooper et al., 2024; Bird and Yibarbuk, 2024; Bird, 2020a; Bettinson and Bird, 2021). If a project damages community trust and harms relationships, not only will that project become endangered, but future projects with the community may also fail due to lack of community support.
Risk 6: Project fails to partner with community where appropriate.
A lack of community participation, generally, can not only lead to poorer data and models, but can also miss opportunities to address social injustices (Sloane et al., 2022). Failure to partner when appropriate with Indigenous communities, specifically, can reproduce histories of ignoring or disrespecting Indigenous sovereignty. Calls for collaborative and participatory research recognise community rights to be meaningfully included in language research that involves them (Carew et al., 2015; Cameron et al., 1993). This is emphasised by unesco in the Los Pinos Declaration calling for a decade of action on Indigenous languages: “Nothing about us without us” (UNESCO, United Nations Educational and Organization).
Risk 7: Project uses culturally unsafe methods to collect data.
Recording one’s own voice can be an uncomfortable experience for many people, and especially so for members of minoritised communities encountering researchers from universities or technology companies using mainstream English. Unless care is taken, prompts might not be culturally relevant, verge into territories of secret traditional knowledge, elicit personal narratives of trauma, or trigger feelings that their culture is not important and valid.
Risk 8: Project does not respect Indigenous data sovereignty.
Failure to adopt appropriate Indigenous data governance protocols can violate the principles of Indigenous people controlling Indigenous data (Carroll et al., 2020; Maiam nayri Wingara, 2018; Te Mana Raraunga, 2016). This is an important aspect of generally recognising Indigenous sovereignty, which was never ceded in what we now know as Australia.
Risk 9: Project does not acknowledge Indigenous knowledge and expertise.
There are many ways in which Indigenous contributions may fail to be acknowledged. Within technology research, this includes not citing or including as co-authors Indigenous experts whose ways of working fall outside Western academic norms (Indigenous Archives Collective et al., 2023).
4. Designing ASR for Australian Aboriginal English
We now discuss a real-world partnership between company anonymized for peer review and university anonymized for peer review which aims to improve ASR for Australian Aboriginal English. We focus on the design considerations, and specifically on our efforts to relate the project goals and processes to both the sociolinguistic context (§2.1) and the technological opportunities (§3.1). We also discuss how we attempt to mitigate the system and project risks (§3.2, §3.4) while providing societal and economic opportunities (§3.3), e.g., by integrating participatory processes throughout the project lifecycle.
When building technologies for Indigenous speech communities—of either ancestral or contact varieties—it is important to consider the ways in which projects face social and technical questions distinct from those that involve mainstream languages (see, e.g., Pine et al., 2024; Bird, 2020b; Carew et al., 2015; Noe and Kirshenbaum, 2024; Lent et al., 2022; Le Ferrand et al., 2022; Taylor et al., 2020). For example, ASR development pipelines typically make assumptions about speech communities and language use, including the existence of ‘gold standard’ spoken and written forms (Markl et al., 2023). These assumptions break down in the case of Aboriginal English, for which the written form is not standardised, and so community participation in the design process is a prerequisite for building technologies that serve the community.
4.1. System desiderata
ASR should support the use of Aboriginal English in all application horizontals which incorporate ASR systems. These include: 1) dictation of messages and documents; 2) voice queries to search engines; 3) voice commands to virtual assistants; and 4) automated transcriptions of voicemail messages, meetings, podcasts, and video content (Aksënova et al., 2021). It is important to distinguish applications in which the transcripts will be processed by computers (Opportunity 4 in §3, e.g., commands to virtual assistants) from those where the transcripts will be read by humans (Opportunities 1–3, e.g., voicemail transcriptions). In the first case, post-processing converts the transcripts to a structured or semi-structured representation, such as an expanded Information Retrieval query or a task frame (Jurafsky and Martin, 2008). This post-processing involves increasing abstraction and standardisation, mapping diverse user inputs into more computationally tractable forms. In the case of transcripts for humans, however, standardisation of linguistic forms is no longer a computational requirement. Instead, standardisation of transcripts minimises the diverse ways in which social meanings are conveyed through linguistic choices. In the process, standardisation reproduces the minoritisation of non-mainstream sociolinguistic identities (Markl et al., 2023). This is especially relevant for informal registers, such as those used in personal messages or social media, in which expressions of social identities are more prevalent and varied.
4.2. Elicitation of speech data
The collection protocols to secure the speech data needed for this project were designed by the authors with advice from the Indigenous Advisory Committee (§4.5). The project also applies participatory research methods discussed further in §7 and previously used by Author 2 and Author 3 (Anonymous, YYYY). However, the data to be collected for this project will be used by company anonymized for peer review and it is expected that the corpus will be open access, in accordance to FAIR and CARE principles (Wilkinson et al., 2016; Carroll et al., 2020), subject to further input from the Indigenous Advisory Committee. For this reason, the project explicitly excludes sharing of taboo cultural materials and sensitive information that is potentially harmful for Indigenous people and communities. To ensure cultural safety (cf. Risk 7), we designed a program of careful elicitation by using culturally relevant and sensitive prompts in consultation with the Indigenous Advisory Committee for this project, and ultimately designed by an Indigenous-owned company which specialises in producing educational resources using an anti-racism lens (Appendix A, cf. Risk 4).
The design of the recording process enables voice data providers to record their voices in their own homes, which is a more comfortable and culturally-sensitive experience than visiting a university or a corporate office. Participant payment processes were designed with both safety and convenience in mind (Opportunity 6).
4.3. Transcription of speech data
The introduction of writing by settler-colonialists has significantly impacted Indigenous populations (Ong, 2013; Franchetto, 2008). In Aboriginal English, there are two major questions regarding orthographic practices. First, how are loan words from ancestral languages, including place names and Country names (e.g. ‘Nyungar’), written in Latin-based script? Second, how do Indigenous people in Australia deliberately adopt their own spelling conventions, e.g., ‘Blak’. These are not purely technical questions, but also sociopolitical (Franchetto, 2008; Markl et al., 2023; Bucholtz, 2000) (cf. Risk 4).
To co-develop transcription guidelines with community members, we required examples of public Aboriginal English to refer to. For this, we used a corpus of over 100 hours of YouTube videos that are primarily in Aboriginal English (see Appendix B). An Indigenous team, led by Author 2, developed the corpus and used their expertise with Aboriginal English to validate the variety spoken, ensuring that speakers were only coded as Aboriginal English if they used at least two linguistic features that distinguish this variety from Standardised Australian English. We developed an initial version of variety-specific transcription guidelines using examples from this corpus. Example guidelines include:
1) borrowings from Pidgin and ancestral languages should be transcribed verbatim, with an English translation included;
2) variety-specific utterance tags should be included, e.g., ‘unna’ and ‘inni’.
The transcription guidelines are a living document, which will grow—with input from the community—as novel classes of examples are observed.
4.4. Evaluation considerations
Doğruöz and Sitaram (2022) have observed that NLP metrics are usually designed with high-resource scenarios in mind. In the case of ASR, the most common evaluation metric is Word Error Rate (WER), which does not actually compare words but rather spelled forms (Karita et al., 2023). It does so in binary fashion: small differences between predictions and reference transcriptions (e.g., ‘naïve’ vs ‘naive’) are penalised just as much as large ones. As such, WER is less useful as a measure of system utility in various contexts, including with language varieties without standardised written forms (Ali et al., 2015; Nigmatulina et al., 2020), and varieties with multiple standard written forms (Kummervold et al., 2024; Karita et al., 2023). In the case of Aboriginal English, WER is not robust to culturally acceptable variation in borrowings from ancestral languages (e.g., ‘Gadigal’ vs ‘Cadigal’ [^4] The Country and people of the Sydney region.) or indigenised spelling variants of English words (e.g., ‘Blak’ vs ‘Black’). [^5] Character Error Rate (CER), another common metric, does not adequately resolve the issue. CER can avoid the overpenalisation of variants accepted by different groups in the community, e.g., only penalising lightly a prediction of ‘Cadigal’ when the reference transcription is ‘Gadigal’ (both refer to the same First Nations Country). However CER is too lenient a metric in cases where the ASR system confuses the spoken word for another word which has a very different meaning or referent, e.g., predicting ‘Kukatj’ when then reference transcription is ‘Kukatja’ (the names of two unrelated Aboriginal language varieties with ancestral lands thousands of kilometres apart). CER would judge such an error as equally bad as the ‘Cadigal’ vs ‘Gadigal’ case.
With a lack of standardised orthography for Aboriginal English, it is important to engage with communities, to centre their perspectives, rather than relying solely on automated metrics such as WER (Cunningham et al., 2024; Markl et al., 2023; Bird, 2020b). As such, the design of the evaluation stage includes interviewing Aboriginal community members, to understand their perspectives on how important fidelity to references is compared to other considerations. An Aboriginal English-specific mapping of commonly accepted spelling variations will be incorporated into WER calculations and enable quantitative measures which provide increased validity with community practices (Ali et al., 2017; Nigmatulina et al., 2020).
4.5. Reflections
We now respond to the questions raised at the start of §3, before providing more specific reflections.
Context.
We acknowledge the colonial context in which this project takes place. Aboriginal and Torres Strait Islander people, the traditional owners of Australia, represent 3.8% of the national population (Australian Government et al., 2020), and are minoritised by the non-Indigenous mainstream population, with significant gaps on economic, educational and health indicators, and life expectancy (Dockery and Milsom, 2007; Commonwealth of Australia, 2024; Australian Institute of Health and Welfare, 2023a, b). Most ancestral languages are extinct or endangered, and Aboriginal English is spoken by around 80% of Aboriginal and Torres Strait Islander people. This project focuses on both Aboriginal English varieties spoken in southwest Western Australia, where Author 2 and Author 3 live and work, and varieties of Aboriginal English spoken in the Pilbara (northern Western Australia), Queensland and Alice Springs (Northern Territory) (cf. Risk 2).
Relationships and participation.
The project is a partnership between company anonymized for peer review and university anonymized for peer review. company anonymized for peer review is a technology company, many of whose products incorporate ASR technologies. university anonymized for peer review is an Australian University whose Linguistics department has both expertise in Aboriginal English and existing relationships with Aboriginal communities, including the Nyungar community. Indigenous associates participate in the project, not just as mere providers of voice data, but as partners in the design of elicitation materials and transcription guidelines, management of the voice data collection, and validation and annotation of the collected materials (Opportunity 7). An Indigenous Advisory Committee was also established to provide guidance throughout the project. Investing in these relationships with partners who have earned community trust helps to mitigate risks to relationships with community (cf. Risk 5).
Control.
The project is funded by company anonymized for peer review, who set the technological goal of the project: improving ASR for Aboriginal and Torres Strait people in Australia.
Appropriateness.
Prior work by Author 1 and Author 4 has identified that improving ASR for Aboriginal English would bring practical benefits to Indigenous people in Australia (Anonymous, YYYY), by supporting spoken language in many applications (cf. Risk 1).
We recognise that the legacy of unethical and settler-colonial research has contributed to distrust of research in many Indigenous communities (Smith, 2021). This, along with a distrust of technology companies, can jeopardise even well intentioned projects. company anonymized for peer review found that seeking out partnership with (socio)linguists at university anonymized for peer review with existing relationships, cultural competencies, and sociolinguistic expertise, was critical, and preferable to attempting to build them from scratch (Doğruöz and Sitaram, 2022; Cooper et al., 2024).
Project members took time to reflect on power dynamics, both within the project and in its context (Noe and Kirshenbaum, 2024; Blodgett et al., 2020) (Opportunity 5). We shared and discussed critiques of extractive data practices. We discussed social injustices, including barriers to opportunity for Indigenous people in Australia, and stigmatisation of Aboriginal English. We strengthened accountability through explicit incorporation of ICIP considerations in our research partnership agreement (§7) (cf. Risk 8).
This project aims to go beyond the baseline of ethical language research, and pursue two higher aspirations (Cameron et al., 1993). 1) Advocacy: we treat Aboriginal English as not merely a topic of research but rather aim to support the speaker community (cf. Schwartz, 2022, sec 3.2). This led us to focus on improving ASR, with tangible benefits. 2) Empowerment: we work with the community, using processes that benefit the community independently of project outputs (cf. Cooper et al., 2024; Cunningham et al., 2024).
It was important from the outset to be clear about which Indigenous people and organisations will participate in the project, and how their participation will be facilitated (Schwartz, 2022; Zhang et al., 2022; Lent et al., 2022; Bird, 2020a; Flavelle and Lachler, 2023; Cooper et al., 2024) (cf. Risk 6). In doing so, it was important to be cognisant of the barriers to technology and research careers facing Indigenous participants, to recognise qualifications and authority beyond that of Western institutions, and to recognise contributors as co-authors of research artefacts (Woods, 2023, p. 219; Meakins et al., 2018, pp. 28–29) (cf. Risk 9). While Indigenous participation is critical to the project, we acknowledge that this does not guarantee that the project will be equitable (Noe and Kirshenbaum, 2024; Sloane et al., 2022). We also recognise that the value given to participation and relationships is asymmetric between various project participants. We were also mindful not to force control onto community partners, who are often under-resourced and overburdened (Cameron et al., 1993; Corbett et al., 2023). Establishing the Indigenous Advisory Committee was an important step in helping to negotiate these tensions. We also incorporated aspects of community knowledge and practices into the design process (Winschiers-Theophilus and Bidwell, 2013) (cf. Risk 6), including using an Indigenous-owned consultancy to design culturally appropriate prompts (cf. Risk 7).
4.6. Summary of mitigation of risks and enabling factors of opportunities
For convenience, in Appendix C we re-iterate and summarise aspects of the projects which mitigate the risks and facilitate the opportunities that we enumerated in §3.
4.7. Limitations
The technological goals of this project were set by company anonymized for peer review, and this precluded Indigenous people exercising full control over the project goals. We also acknowledge that Aboriginal and Torres Strait Islander people are less likely to be in full time research positions at companies or universities in Australia, limiting opportunities to be partner investigators on this research.
We acknowledge the plurality of Indigenous cultures and Countries in Australia, and that Indigenous co-investigators, designers, managers and transcribers cannot represent the full diversity of those cultures and Countries. Our decision to include different regions is important not only because of lexico-grammatical differences, but also because Aboriginal English has a key emblematic function as an encoder of different ways of being and doing. We believe that the various ancestries represented wherever Aboriginal English is used are as important as understanding Aboriginal English itself.
Our corpus for developing the initial version of the transcription guidelines was limited to publicly shared YouTube videos. We do not claim that it is representative of spoken Aboriginal English in other domains, such as personal messages, nor of written forms of Aboriginal English.
As researchers based in Australia, our lived experiences with colonial contexts are also based primarily in Australia. Although we hope that this project contains aspects that generalise, we do not claim to know or represent other contexts.
5. Discussion: Further Technology Opportunities for Australian Aboriginal English
We now consider technologies supporting Aboriginal English speakers beyond ASR. Given that languages go by multiple names, the NLP community typically uses language codes as identifiers in multilingual systems, datasets, and data catalogues (Kreutzer et al., 2022). This discrete mapping of languages to codes is central to NLP’s conceptual model of what languages are (Smart et al., 2024). Aboriginal English currently lacks a standardised identifier, and this would be useful for future work on this variety. However, given the power of classifying (Bowker and Star, 2000)—and the colonial associations of the origins of the ISO 639-3 identifiers (Dobrin and Good, 2009), which also underpin codes such as the BCP 47—we echo calls for involving community members and linguists, and avoiding offensive exonyms, when designing codes for Aboriginal English (Morey et al., 2013).
Given the primary orality of Aboriginal English, and its distinct grammatical and discourse-pragmatic features relevant to information seeking (§2.1), there are open questions around whether spoken language assistants could better support speakers. Robust support for tag-questions—rather than just questions with subject-verb inversion—might provide more natural support for information requests. While speakers of Aboriginal English benefit from hearing others talk in the same way (Angelo, 2021, p. 42), the question of whether digital assistants should try to ‘sound Indigenous,’ must be approached with care, to avoid potentially reproducing interactions of racialised subservience to non-Indigenous consumers (Owens, 2023).
Before developing writing tools tailored for Aboriginal English, more understanding is required of how writing practices diverge from standardised Australian English. Social media and other domains with informal registers are more likely to have written Aboriginal English, and it may be that modeling orthographic variation is a useful approach for text technologies (cf.  Lin et al., 2024).
Evidence suggests that designing health campaigns in Aboriginal English has community benefits (Collard and Rodríguez Louro, 2025). This suggests potential use cases for machine translation, particularly automated translation and rewriting of health and other social services materials in Aboriginal English.
Finally, given the prevalence of code switching between Aboriginal English and Standardised Australian English, technologies which support the former would benefit from also supporting the latter. Technologies which aim to support English in Australia should explicitly document whether they support Aboriginal English (Bender, 2019).
6. Conclusions
This paper has explored three questions. First, can speech technology projects support speakers of Australian Aboriginal English by affording specific opportunities? Second, what are the system deployment and procedural risks of such endeavours? Third, how can we mitigate risks, and facilitate opportunities, by integrating participatory practices which appropriately respect the context and are culturally appropriate for the community? We explored these questions through a detailed study of the design considerations of a real-world project that aims to improve ASR for Aboriginal English. The case study of ASR also brings greater attention to a more general trend: the lack of technological support for Indigenous contact language varieties, and the ways that this perpetuates neo-colonialism through a continued failure to recognise and support Indigenous socio-cultural identities and ways of being. Contact language varieties are critical to contemporary Indigenous sociolinguistic identities, and to civil and economic participation, including in places such as Australia where settler-colonialism has led to the extinction of most ancestral languages. We discussed a number of ways in which speech technologies can support speakers of Aboriginal English, as well as the risks of harms caused by system errors. This paper contributes to our understanding of ethical and responsible development of technologies for minoritised language communities, as well as to our understanding of practices for building culturally sensitive technologies while incorporating participatory methods. Although the project focuses on a specific technology, community and context—ASR for speakers of an indigenised variety of English in what we now know as Australia—we believe that some of the considerations and methodologies may be applicable to other technology projects.
7. Ethical considerations
Before funding was secured for the project, Aboriginal scholars (with expertise in education, applied linguistics, and language revitalisation) were approached and asked to participate as members of a committee advising researchers on the project. Because the research involves Indigenous participants, letters of support from these advisors and from aiatsis (Australian Institute of Aboriginal and Torres Strait Islander Studies) were included in a submission to university anonymized for peer review’s IRB. Both university anonymized for peer review’s IRB and company anonymized for peer review’s privacy and legal teams reviewed the consent forms that would be used with Aboriginal people creating voice recordings.
In addition, we sought and obtained the support and expertise of the Language Data Commons of Australia (LDaCA) [^6] https://www.ldaca.edu.au/, as the research involves creating a new dataset of Indigenous language materials. LDaCA advised on culturally appropriate data governance and access frameworks (Foley et al., 2024).
The collection of Aboriginal English voice recordings for the purposes of this project raises ICIP issues. As a team, the academic authors at university anonymized for peer review negotiated with company anonymized for peer review the inclusion of an ICIP clause in the research agreement signed in November 2023. This clause reads as follows:
The parties acknowledge that Australian protocols may be relevant to ICIP in relation to the Research and will work collaboratively during the Research Term to agree the extent to which Australian protocols, including but not limited to the Australian Council for the Arts Protocols for using First Nations Cultural and Intellectual Property in the Arts (2019), should apply to the Research.
The collection of Aboriginal English for use by company anonymized for peer review must be approached with extreme care. From the start, Author 2 and Author 3 clarified with their partners that they would not use unstructured yarning sessions like they had done for Anonymous Grant Name (see, e.g., (Anonymous, YYYY)). With no pre-determined questions and clear Indigenous leadership through the research process, that project collected language as used in everyday settings, including sensitive stories about dispossession, trauma and death by suicide. These materials are firmly closed access with no risk of the audio files and associated transcriptions entering the public domain. However, unstructured yarning sessions tend to give rise to highly culturally sensitive material, which should be protected from use by third parties, especially large technology companies like company anonymized for peer review. For this reason, in consultation with Author 1 and the Advisory Committee, Author 2 and Author 3 recommended for this project the design of conversational prompts and story boards (see Appendix A) which seek to elicit interaction on a variety of general topics while steering clear of culturally sensitive material.
Once collected, all data will be screened by a team of Indigenous RAs, led by Author 3, and consisting of anonymised for peer review and anonymised for peer review, who will vet the audio files and transcriptions for cultural sensitivity. Any culturally unsafe material (e.g., stories of specific people, places, Indigenous knowledges, trauma, and loss) will be erased from the corpus before sharing the data with others.
Appendix A: Visual prompts for eliciting speech data collection
The prompts in Figure 1 were designed by Indigenous-owned consultancy Goorlil Consulting.
(a) Food
(b) Fishing
(c) Sport
(d) Visiting family
(e) Hanging out at the shops
(f) Fun times and funny jokes
(g) Motorcars
Figure 1. Visual prompts for speech data elicitation.
Appendix B: Composition of corpus of Aboriginal English on YouTube
For readers who wish to hear examples of Aboriginal English, we recommend searching on YouTube using queries such as “My art, My culture”, “Telling Our Stories, Our Stolen Generations”, “Through our Eyes, Series 3”, or “Community Spotlight Artists Profile Cairns Indigenous arts fair”.
YouTube grants users of the platform ‘a worldwide, non-exclusive, royalty-free license to access your Content through the Service, and to use that Content, including to reproduce, distribute, prepare derivative works, display, and perform it’ (https://www.youtube.com/static?template=terms). Our use of YouTube videos to create transcription guidelines was consistent with this. These transcription guidelines do not contain any personal identifying information of offensive content.
The YouTube video corpus that we describe in §4 has the composition summarised in Tables 1–3.
Domain # videos Total
duration
Arts and Culture 203 23.63
Autos and Vehicles 4 0.47
Comedy 29 2.20
Education 90 14.38
Entertainment 16 2.60
Film and Animation 30 7.82
Gaming 0 0
Health 26 2.53
How to / Style 14 1.79
News and Politics 42 6.25
Nonprofits and Activism 102 14.31
People and Blogs 149 22.29
Pets and Animals 7 0.82
Science and Technology 25 4.00
Sports 17 1.91
Travel and Events 22 4.37
Total 776 111.37
Table 1. Domains of videos.
Scripted? # videos Total duration (hours)
Yes 134 18.31
No 510 77.17
Unknown 132 15.89
Total 776 111.37
Table 2. Whether videos were scripted.
Percent in # videos Total duration
Aboriginal English (hours)
[50, 60) 34 8.43
[60, 70) 34 5.77
[70, 100] 708 97.18
Total 776 111.37
Table 3. Percent of speech in Aboriginal English in videos. Only videos with at least 50% were considered; 70+% was preferred.
Appendix C: Summaries of mitigations of risks and enabling factors of opportunities
Risk Mitigated by
1. Deploying inappropriate technology Building on prior work that identified the appropriateness of ASR for this context; Validation through expert partners with expertise working with this community.
2. Low ecological validity Multi-region data collection to guarantee that regional variations are represented; In-person data collection techniques led by Indigenous employees, so that speakers are comfortable speaking naturally.
3. Psychological harms from ASR errors Primary goal of the project: reducing ASR errors for Aboriginal English. We also call for more HCI research both on potential harms experienced by Indigenous communities, as well as on potential mitigations in the application layer.
4. Torquing identities Community involvement in transcription guidelines; We call for more work on the transcription preferences of speakers of non-standardised languages.
5. Harming relationships Partnering with experts with existing relationships with community; Establishing an Indigenous Advisory Committee.
6. Failing to partner with community Indigenous researcher on the project; Hiring Indigenous staff; Engaging an Indigenous-owned consultancy; Establishing an Indigenous Advisory Committee.
7. Culturally unsafe methods in data collection In-person, Indigenous-led data collection; Using culturally safe prompts developed by an Indigenous-owned consultancy; Verifying collected data for cultural safety; Researchers on the project with experience running culturally safe data collections.
8. Not respecting Indigenous data sovereignty company anonymized for peer review not owning the collected speech data; Indigenous Advisory Committee guidance on appropriate Indigenous Data Governance.
9. Not acknowledging Indigenous knowledge and expertise Recognising contributors as co-authors on research artefacts; Engaging an Indigenous-owned consultancy; Establishing an Indigenous Advisory Committee.
Table 4. Project Risks and Mitigations
Opportunity Enabled by
1–4. Deployed system helpful to community and non-community members in diverse application contexts Explicitly including support for various application contexts in design desiderata.
5. Project members develop greater understanding of Indigenous contexts and histories Team discussions on power dynamics, barriers, and social injustices; Team discussions about language use during the writing of this paper and other artefacts.
6. Economic opportunities Equitable payment for Indigenous data providers; Employing Indigenous staff.
7. Building capabilities Employing Indigenous staff in roles that enable skill development, including designing elicitation materials and transcription guidelines, and managing voice data collection.
8. Self-determination Pursuing data governance frameworks which give Indigenous people the power to decide how the Indigenous voice recordings might be used in the future.
Table 5. Project Opportunities and Enabling Factors